OpenStudy (anonymous):

What is the percent of increase from 50 to 90? A. 40% B. 44% C. 56% D. 80%

TheSmartOne (thesmartone):

So what is 90-50=? @ohva.student

OpenStudy (anonymous):

Though that's not all the problem is, 90-50 = 40 but then you have to divide by 50 to see the percent increase.

OpenStudy (anonymous):

will medal and fan

TheSmartOne (thesmartone):

@ohva.student First what is 90-50=? Then we can get to the next step to arrive to your answer :)

OpenStudy (anonymous):

40 @TheSmartOne

TheSmartOne (thesmartone):

So what is 40/50=?

OpenStudy (anonymous):

0.8

OpenStudy (anonymous):

x100 that's your answer

TheSmartOne (thesmartone):

^ :D

OpenStudy (anonymous):

80

TheSmartOne (thesmartone):

Don't forget the % sign :P So its actually 80% not just 80 :D

OpenStudy (anonymous):

ok thanks so must @iambatman and @TheSmartOne